ExplodingDryer:

A triangle has sides that measure 2 units, 5 units, and 5.39 units. What is the area of a circle with a circumference that equals the perimeter of the triangle? Use 3.14 for π, and round your answer to the nearest whole number.

ExplodingDryer:

plz help

TheSmartOne:

Perimeter of a triangle = sum of a sides What would the perimeter be? Add up those three numbers and that's your perimeter. Now, circumference = \( 2\pi r\) plug in that number you just got for the perimeter with the circumference solve for r So basically divide by \( 2 \pi\) on both sides You should now have r = some number Finally, to find the area Area = \(\pi r^2\) Plug in that value for r that you just got and calculate the area And that would be the area of a circle that has a circumference that is equal to the perimeter of that triangle

TheSmartOne:

Let me know what you get or if you'd like further clarification! I'd be more than happy to help :)